# Subdiffusive transport in the Fredkin dynamical universality class

### v4

@@ II.1 The Fredkin constraint
The most general formulation of the Fredkin constraint involves mapping a system’s local degrees of freedom to a height field. For example, a binary system in one dimension may be interpreted as a 1D lattice consisting of particles and holes. Mapping this system to a height field requires starting at some point $x=0$ where $h(0)=0$; the height field at any other point is given by $h(x)=\sum_{n<x}\Delta h_{n}$, where $\Delta h_{n}=1$ for a particle and $\Delta h_{n}=-1$ for a hole (fig. 1a). The Fredkin constraint forbids any rearrangement of local degrees of freedom that decreases the local minimum value of the height field. In the particle/ hole language, the height field constraint can be interpreted as a 4-site “controlled-swap” restriction on local rearrangements of charges (fig. 1b).
II.2 The Fredkin spin chain
While Fredkin-constrained dynamics are generic and may appear in many contexts, the derivation of the dynamical exponent in this work will make use of the correspondence between the dynamics of Markov processes and the gap scaling of critical frustration-free Hamiltonians at a Rokshar-Kivelson point U1FRUC ; SanjayAmosFracton ; hans_subdiffusion . We will therefore focus on a pair of one-dimensional quantum spin chains that map to Fredkin-constrained Markov processes: the Fredkin (spin-$\frac{1}{2}$)  Korepin2016fredkin and Motzkin (spin-1)PhysRevLett.109.207202 ; FredkinEntanglement chains. Both spin chain variants have frustration-free Hamiltonians with an exactly-known critical ground state. The energy gap of these Hamiltonians are rigorously upper-bounded as $L^{-2}$ for a system of size $L$ movassagh2016gap —therefore, despite being a one-dimensional critical point, they are not described by a conformal field theory ChenFredkin2017 ; ChenFredkin2017a .
For brevity, we will limit the discussion in the main text to the Fredkin spin chain, and all results are reproduced for the Motzkin spin chain in Appendix D. The Fredkin Hamiltonian is given by
@@ II.2 The Fredkin spin chain
Figure 1: The matched parentheses charge and the Fredkin constraint (a) Three representations (height field, random walkers, and Dyck alphabet) of a classical Markov process in one dimension with two degrees of freedom. (b) The Fredkin constraint prohibits rearrangements of local degrees of freedom that decrease the local minimum value of the height field. In each configuration shown, the particle may hop between the second and third sites if the neighboring sites satisfy the height-field constraint; the three moves that satisfy the height field constraint are shown on the top, while the disallowed move that changes the local minimum of the height field is shown at the bottom. In the Dyck alphabet representation, the allowed moves locally conserve the number (but not the type) of irreducible Dyck words, while the disallowed move destroys a Dyck word.
II.3 Dyck words
As originally outlined in Korepin2016fredkin , the Fredkin Hamiltonian naturally lends itself to a combinatorial description in terms of the Dyck alphabet. We briefly summarize some important aspects of the Dyck alphabet here. The Dyck alphabet consists of two symbols: open and closed parentheses. Dyck words are defined as strings of parentheses in which every parenthesis is matched when read from left to right; for instance, (()) and ()() are both Dyck words, while )()( is not. A Dyck word necessarily contains an even number of parentheses and can be identified by its semilength, $\ell$, which is half of the total number of parentheses. The number of Dyck words of semilength $\ell$ is given by the $\ell^{\rm th}$ Catalan number, $C_{\ell}$. A class of Dyck words that will be important for the following discussion are irreducible Dyck words. Irreducible Dyck words are those in which the first parenthesis in the word is matched by the final parenthesis in the word; for example, ((())) and (()()) are both irreducible Dyck words of semilength $\ell=3$, while (())(), ()(()), and ()()() are also Dyck words of semilength $\ell=3$ but are not irreducible. Since irreducible Dyck words of semilength $\ell$ are of the form (…), where … must itself be a Dyck word, there are $C_{\ell-1}$ irreducible words of semilength $\ell$.
@@ III The matched parentheses charge
To illustrate the conservation law, consider the configuration (()(()))(()(()))( ( ) ( ( ) ) ). There are two $\ell=1$ words in this configuration in addition to two other irreducible words of semilengths $\ell=2$ and $4$; each of these four irreducible words corresponds to a single matched parentheses charge. The position of the matched parentheses charge is defined as the midpoint of the underlying irreducible Dyck word, which lives on a link– for example, the $\ell=2$ word in the above configuration has position $n=5$. We label the matched parentheses charges as $q_{n}^{\ell}$ and adopt the convention that the lower index denotes position and the upper index denotes semilength; if either index is irrelevant to the discussion at hand, it will be omitted.
In the configuration (()(()))(()(()))( ( ) ( ( ) ) ), Fredkin dynamics permits swapping the second and third parentheses, leading to the move: $\underline{(\underline{(|)}(|()))}\rightarrow\underline{(|)}~{}\underline{(((|)))}$. The charges affected by this move (initially $q_{2}^{1}$ and $q_{4}^{4}$) are underlined and their positions are marked with a vertical bar. The illustrated move consists of two distinct parts: the movement of the initial $q_{2}^{1}$ charge to the left, and the response of the original $q_{4}^{4}$ charge, which changes semilength from $\ell=4$ to $\ell=3$ as it moves to the right.
To generalize the above discussion, moves permitted by the Fredkin Markov process can be decomposed into two parts: the movement of a single $q^{1}$ charge and the response of some other $q^{\ell}$ charge. As illustrated in the inset of Fig. 1(b), allowed moves correspond to moving a $q^{1}$ charge to either the left or right by a single link. At $S^{z}_{\text{tot}}=0$, a $q^{1}$ charge necessarily passes through the end parenthesis of some other $q^{\ell}$ charge, which in turn causes the $q^{\ell}$ charge to both move by one site and change in semilength to ±$\ell\pm 1$. These processes are the only moves allowed by the Fredkin Markov process.
In the spin chain language, the Fredkin Hamiltonian may be rewritten in terms of the matched parentheses charge’s hard-core bosonic creation and annihilation operators; each term exactly corresponds to the movement of a $q^{1}$ charge in the Markov process. Writing $(a_{i}^{\ell})^{\dagger}$ as the creation operator for a matched parentheses charge centered on link $i$ with semilength $\ell$, the Hamiltonian may be expressed as:
$\begin{split}&H=\frac{1}{2}\sum_{i,\ell}n_{i-1}^{\ell+1}n_{i-\ell}^{1}+n_{i}^{\ell}n_{i-\ell-1}^{1}+n_{i+1}^{\ell+1}n_{i+\ell}^{1}+n_{i}^{\ell}n_{i+\ell+1}^{1}\\ &-(a_{i-1}^{\ell+1})^{\dagger}(a_{i-\ell}^{1})^{\dagger}a_{i}^{\ell}a_{i-\ell-1}^{1}-(a_{i+1}^{\ell+1})^{\dagger}(a_{i+\ell}^{1})^{\dagger}a_{i}^{\ell}a_{i+\ell+1}^{1}-\text{h.c.},\end{split}$ (2)
where $n_{i}^{\ell}=(a_{i}^{\ell})^{\dagger}a_{i}^{\ell}$. In this representation one can see that the total matched parentheses charge is $Q_{\rm tot}=\sum_{i,\ell}n_{i}^{\ell}$.
IV Pseudolocality of the matched parentheses chargr
@@
The total matched parentheses charge $Q_{\rm tot}=\sum_{i=1}^{L}q_{i}$ is related to $S^{z}_{\rm tot}$; for example, $Q_{\rm tot}=L/2$ for a system of $L$ spins with $S^{z}_{\rm tot}=0$ with periodic boundary conditions. As a result, any spin-$\frac{1}{2}$ Hamiltonian that preserves $S^{z}_{\rm tot}$ acting on a system with periodic boundary conditions will globally conserve the number of matched parentheses charges. In contrast, the Fredkin Hamiltonian has a stronger, local conservation law governing the matched parentheses charge and so one has a corresponding continuity equation (see Appendix A). The locality of the matched parentheses charge is not entirely obvious a priori– the total matched parentheses charge may be written as the sum of terms $Q_{\rm tot}=\sum_{n}q_{n}$, but the support of each $q_{n}$ is not obviously bounded. Here, we first show that the matched parentheses charge has power-law tails that account for nontrivial static correlations between parentheses. We then prove that the charge is additive and extensive, thus demonstrating that even with its power-law tails, the charge is relevant to transport.
@@
First, we consider the source of the algebraically-decaying tails of the matched parentheses charge. All static correlations between matched parentheses charges arise from the fact that a parenthesis may only function as the endpoint of a single $q^{\ell}$ charge. To illustrate this, consider a single $q_{0}^{1}$ charge, corresponding to a configuration ()()( ) at the origin. Since parentheses may not be shared between charges, and since the existence of a $q_{\pm 1}^{1}$ charge would require one of the endpoints of the $q_{0}^{1}$ to instead be one of the endpoints of the $q_{\pm 1}^{1}$ charge, the simultaneous existence of both charges is impossible. Extending this logic to an arbitrary link $\ell$, the presence of $q_{0}^{1}$ precludes the existence of any $q_{\pm\ell}^{\ell}$ charge. The tails of the matched parentheses charge therefore depend on the probability of finding a $q^{\ell}$ charge $\ell$ away from a $q^{1}$ charge. Since the probability of finding a $q^{\ell}$ charge on a generic site is given by $C_{\ell-1}/2^{2\ell}\sim\ell^{-3/2}$, all nontrivial static correlations between the presence of $q_{0}^{1}$ and any other charges are given solely by the probability of finding $q_{\pm\ell}^{\ell}$. Therefore, we find $\langle q_{m}^{1}q_{n}\rangle\sim|m-n|^{-3/2}$.
@@
For a charge to be relevant to thermodynamics, it must be extensive and additive; despite having power-law tails, the matched parentheses charge $q^{\ell}$ indeed exhibits both of these properties. To see this, consider a region of $2L\rightarrow\infty$ total sites split into two $L$-site subsystems $A$ and $B$. The total charge in this region $Q_{AB}=\sum_{n=1}^{2L}q_{n}$. The total number of matched parentheses charges in subsystem $A$, given by $Q_{A}=\sum_{n=1}^{L}q_{n}$, grows with $L$, and so $\langle Q_{A}\rangle\sim L$. One could worry that in spite of the extensivity of the matched parentheses charge, the power-law tails of the charge may lead to $Q_{AB}\neq Q_{A}+Q_{B}$. From showing exactly that $\langle Q_{A}^{2}\rangle_{c}\sim L$, we conclude that fluctuations of the total charge are sufficiently suppressed to guarantee additivity $Q_{AB}\simeq Q_{A}+Q_{B}$, up to subleading corrections of $\mathcal{O}(\sqrt{L})$ – for further details, see Appendix A. Therefore, fluctuations in the value of $Q_{A}$ are suppressed as $\sqrt{\langle Q_{A}^{2}\rangle_{c}}/{\langle Q_{A}\rangle}\sim 1/\sqrt{L}$, so the local density of matched parentheses is well-defined. Since the matched parentheses charge is locally conserved, extensive, and additive, it is relevant to hydrodynamics. Intuitively, due to the fact that each matched parentheses charge is tied to two individual parentheses, both of which move with $z\approx 5/2$, the matched parentheses charge should have the same dynamical exponent as the original charge– this intuition can be confirmed numerically (see Appendix C).

### v5

@@ II.1 The Fredkin constraint
The most general formulation of the Fredkin constraint involves mapping a system’s local degrees of freedom to a height field. For example, a binary system in one dimension may be interpreted as a 1D lattice consisting of particles and holes. Mapping this system to a height field requires starting at some point $x=0$ where $h(0)=0$; the height field at any other point is given by $h(x)=\sum_{n<x}\Delta h_{n}$, where $\Delta h_{n}=1$ for a particle and $\Delta h_{n}=-1$ for a hole (fig. 1a). The Fredkin constraint forbids any rearrangement of local degrees of freedom that decreases the local minimum value of the height field. In the particle/ hole language, the height field constraint can be interpreted as a 4-site “controlled-swap” restriction on local rearrangements of charges (fig. 1b) although in principle, the height-field constraint may be implemented over an arbitrary number of sites.
II.2 Mapping between Markov processes and quantum spin chains
While Fredkin-constrained dynamics are generic and may appear in many contexts, the derivation of the dynamical exponent in this work will make use of a correspondence between the dynamics of Markov processes and the gap scaling of critical frustration-free Hamiltonians at a Rokshar-Kivelson (RK) point U1FRUC ; SanjayAmosFracton ; hans_subdiffusion .
Consider a classical discrete-time Markov process described by a probability density $\vec{p}(t)$ over all possible states. At time $t=0$, let $p_{\alpha}(0)=1$ and all $p_{\beta\neq\alpha}(0)=0$. Under stochastic time evolution, the probability density will relax to an equilibrium state in which all possible states are assigned equal probabilities. The mixing time associated with this relaxation process can be expressed as the inverse of the “gap” of this Markov process, $\Delta_{M}$; if the Markov process is gapless, we expect $\Delta\rightarrow 0$ in the thermodynamic limit.
The discrete-time Markov process can be viewed as a Trotterized version of a stochastic continuous-time process $\vec{p}(t+\tau)=T\vec{p}(t)$, where $T$ is the classical transfer matrix describing the Markov process. In the continuous time limit $\tau\to 0$, the transfer matrix $T$ may be related to a quantum Hamiltonian $H$ through the usual relation $T\simeq 1-\tau H$. The class of quantum Hamiltonians that are related to Markov processes through standard quantum-classical mappings are known as Rokhsar-Kivelson (RK, or frustration-free) Hamiltonians. The gap of the Markov matrix $\Delta_{M}$ is then directly related to the gap $\Delta$ of $H$, and therefore, the scaling of $\Delta_{M}$ is the same as the gap-scaling of the RK Hamiltonian $\Delta_{M}\sim\Delta\sim L^{-z}$ SanjayAmosFracton .
II.3 The Fredkin spin chain
To make use of the mapping described in the previous section, in this work we focus on a pair of one-dimensional quantum spin chains that map to Fredkin-constrained Markov processes: the Fredkin (spin-$\frac{1}{2}$)  Korepin2016fredkin and Motzkin (spin-1)PhysRevLett.109.207202 ; FredkinEntanglement chains. Both spin chain variants have frustration-free Hamiltonians with an exactly-known critical ground state. The energy gap of these Hamiltonians are rigorously upper-bounded as $L^{-2}$ for a system of size $L$ movassagh2016gap —therefore, despite being a one-dimensional critical point, they are not described by a conformal field theory ChenFredkin2017 ; ChenFredkin2017a .
For brevity, we limit the discussion in the main text to the Fredkin spin chain, and all results are reproduced for the Motzkin spin chain in Appendix D. The Fredkin Hamiltonian is given by
@@ II.3 The Fredkin spin chain
Figure 1: The matched parentheses charge and the Fredkin constraint (a) Three representations (height field, random walkers, and Dyck alphabet) of a classical Markov process in one dimension with two degrees of freedom. (b) The Fredkin constraint prohibits rearrangements of local degrees of freedom that decrease the local minimum value of the height field. In each configuration shown, the particle may hop between the second and third sites if the neighboring sites satisfy the height-field constraint; the three moves that satisfy the height field constraint are shown on the top, while the disallowed move that changes the local minimum of the height field is shown at the bottom. In the Dyck alphabet representation, the allowed moves locally conserve the number (but not the type) of irreducible Dyck words, while the disallowed move destroys a Dyck word. (c) A configuration of parentheses (top) translated into a configuration of matched parentheses charges (bottom). Each matched parentheses charge sits on on the central link $n$ between the two matched parentheses and has semilength $\ell$ that is half of the distance between the matched parentheses. (d) An example of a move allowed under the Fredkin Markov process, in which the second and third parentheses in the configuration swap positions. Allowed Fredkin moves conserve the number of matched parentheses charges and correspond to the local hopping of these charges.
II.4 Dyck words
As originally outlined in Korepin2016fredkin , the Fredkin Hamiltonian naturally lends itself to a combinatorial description in terms of the Dyck alphabet. We briefly summarize some important aspects of the Dyck alphabet here. The Dyck alphabet consists of two symbols: open and closed parentheses. Dyck words are defined as strings of parentheses in which every parenthesis is matched when read from left to right; for instance, (()) and ()() are both Dyck words, while )()( is not. A Dyck word necessarily contains an even number of parentheses and can be identified by its semilength, $\ell$, which is half of the total number of parentheses. The number of Dyck words of semilength $\ell$ is given by the $\ell^{\rm th}$ Catalan number, $C_{\ell}$, which scales asymptotically as $C_{\ell}\sim\ell^{3/2}$. A class of Dyck words that will be important for the following discussion are irreducible Dyck words. Irreducible Dyck words are those in which the first parenthesis in the word is matched by the final parenthesis in the word; for example, ((())) and (()()) are both irreducible Dyck words of semilength $\ell=3$, while (())(), ()(()), and ()()() are also Dyck words of semilength $\ell=3$ but are not irreducible. Since irreducible Dyck words of semilength $\ell$ are of the form (…), where … must itself be a Dyck word, there are $C_{\ell-1}$ irreducible words of semilength $\ell$.
@@ III The matched parentheses charge
To illustrate the conservation law, consider the configuration (()()())(()()())( ( ) ( ) ( ) ). This configuration consists of four irreducible Dyck words, three of semilength $\ell=1$ and one of semilength $\ell=4$. We define the position of each irreducible Dyck word by the location of its midpoint; in this configuration, the $\ell=1$ words are located on links $n=2,4,$ and $6$, while the midpoint of the $\ell=4$ word falls on link $n=4$. Consider swapping the second and third sites in the above configuration, which is permitted under Fredkin-constrained dynamics. The word of semilength $\ell=4$ and the $\ell=1$ word originally located on link $n=2$ are affected by this move, while the other two $\ell=1$ words are unaffected. This process is illustrated below, with the affected words underlined with their midpoints marked with a vertical bar:
$\underline{(\underline{(|)}(|)())}\rightarrow\underline{(|)}~{}\underline{(()|())}.$ (2)
After this allowed move, the word that was originally located on link $n=2$ with semilength $\ell=1$ has moved to link $n=1$, while the word that originally had semilength $\ell=4$ whose midpoint was located on link $n=4$ has moved to link $n=5$ and has decreased in semilength to $\ell=3$.
This example captures a general characteristic of moves permitted under Fredkin constraints: all allowed moves can be interpreted as a Dyck word of semilength $\ell=1$ moving past another parenthesis (see fig. 11b). In the $S_{z}^{\rm tot}=0$ charge sector with periodic boundary conditions, every parenthesis is the endpoint of a Dyck word. The effect of moving an $\ell=1$ Dyck word past the endpoint of some other word of semilength $\ell$ is to move that word’s midpoint by a single link to $n\pm 1$ while changing its semilength by one to $\ell\mp 1$.
It is also instructive to consider a move that is forbidden under Fredkin-constrained dynamics:
(()()())↛(())(()).↛(()()()$\not\rightarrow$(())(()).( ( ) ( ) ( ) ) ↛ ( ( ) ) ( ( ) ) . (3)
The initial configuration has two irreducible Dyck words of semilengths $\ell=1$ and $\ell=4$ whose midpoints fall on link $n=4$. The final configuration corresponds to the destruction of both of these Dyck words and the creation of two $\ell=2$ Dyck words whose midpoints fall on links $n=2$ and $n=6$. In contrast to the process depicted in 2, the process shown in 3 does not involve the movement of an $\ell=1$ Dyck word past the endpoint of some other Dyck word. Therefore, despite the fact that globally, the total number of irreducible Dyck words remains unchanged, the process illustrated in 3 violates the local conservation of irreducible Dyck words.
In summary, all allowed Fredkin moves can be interpreted as the local hopping of irreducible Dyck words, while moves forbidden under Fredkin-constrained dynamics correspond to the local destruction (and creation) of Dyck words. The midpoint of an irreducible Dyck word can then be interpreted as a conserved charge– we accordingly call this the “matched parentheses charge.” Each matched parentheses charge has two labels– its position, $n$, and its semilength, $\ell$. In the $S_{z}^{\rm tot}$ charge sector with periodic boundary conditions, the hopping of an $\ell=1$ charge is always accompanied by the movement of another charge– if this charge was originally of semilength $\ell$ and located on link $n$, then moving the $\ell=1$ charge past one of its endpoints will cause it to hop to link $n\pm 1$ and change in semilength to ±$\ell\pm 1$.
The matched parentheses charge is also conserved by the Hamiltonian of the Fredkin quantum spin chain. The Fredkin Hamiltonian may be re-written so as to make the presence of the matched parentheses charge more apparent. First, we introduce the hard-core bosonic creation and annihilation operators of the matched parentheses charge, $(a_{i}^{\ell})^{(\dagger)}$. The number operator may be written as $q_{i}^{\ell}=(a_{i}^{\ell})^{\dagger}a_{i}^{\ell}$. In the discussion that follows, we adopt the convention that the lower index denotes position and the upper index denotes semilength; if either index is irrelevant to the discussion at hand, it will be omitted and assumed to be summed over.
Within the $S_{z}=0$ charge sector with periodic boundary conditions, the Hamiltonian may then be expressed as:
$\begin{split}&H=\frac{1}{2}\sum_{i,\ell}q_{i-1}^{\ell+1}q_{i-\ell}^{1}+q_{i}^{\ell}q_{i-\ell-1}^{1}+q_{i+1}^{\ell+1}q_{i+\ell}^{1}+q_{i}^{\ell}q_{i+\ell+1}^{1}\\ &-(a_{i-1}^{\ell+1})^{\dagger}(a_{i-\ell}^{1})^{\dagger}a_{i}^{\ell}a_{i-\ell-1}^{1}-(a_{i+1}^{\ell+1})^{\dagger}(a_{i+\ell}^{1})^{\dagger}a_{i}^{\ell}a_{i+\ell+1}^{1}-\text{h.c.},\end{split}$ (4)
In this representation one can see that the total matched parentheses charge is $Q_{\rm tot}=\sum_{i,\ell}q_{i}^{\ell}$.
IV Pseudolocality of the matched parentheses charge
@@
The total matched parentheses charge $Q_{\rm tot}=\sum_{n}q_{n}$ is related to $S^{z}_{\rm tot}$; for example, $Q_{\rm tot}=L/2$ for a system of $L$ spins with $S^{z}_{\rm tot}=0$ with periodic boundary conditions. As a result, any spin-$\frac{1}{2}$ Hamiltonian that preserves $S^{z}_{\rm tot}$ acting on a system with periodic boundary conditions will globally conserve the number of matched parentheses charges. In contrast, the Fredkin Hamiltonian has a stronger, local conservation law governing the matched parentheses charge and so one has a corresponding continuity equation (see Appendix A). The locality of the matched parentheses charge is not entirely obvious a priori– the total matched parentheses charge may be written as the sum of terms $Q_{\rm tot}=\sum_{n}q_{n}$, but the support of each charge is not obviously bounded. Here, we first show that the matched parentheses charge has power-law tails that account for nontrivial static correlations between parentheses. We then prove that the charge is additive and extensive, thus demonstrating that even with its power-law tails, the charge is relevant to transport.
@@
First, we consider the source of the algebraically-decaying tails of the matched parentheses charge. All static correlations between matched parentheses charges arise from the fact that a parenthesis may only function as the endpoint of a single charge of semilength $\ell$. To illustrate this, consider a charge corresponding to a configuration ()()( ) at the origin, $n=0$ and $\ell=1$. Since parentheses may not be shared between charges, and since the existence of a charge of semilength $\ell^{\prime}=1$ at site $m=\pm 1$ charge would require one of the endpoints of the $n=0,\ell=1$ charge to instead be one of the endpoints of the $m=\pm 1,\ell^{\prime}=1$ charge, the simultaneous existence of both charges is impossible. Extending this logic to an arbitrary link $\ell^{\prime}$, the presence of a charge with $n=0,\ell=1$ precludes the existence of any $m=\pm\ell^{\prime},\ell^{\prime}$ charge. The tails of the matched parentheses charge therefore depend on the probability of finding a charge of semilength $\ell^{\prime}$ that lives $\ell^{\prime}$ links away from a charge of semilength $\ell=1$. Since the probability of finding a charge of semilength $\ell$ on an arbitrary site is given by $C_{\ell-1}/2^{2\ell}\sim\ell^{-3/2}$, all nontrivial static correlations between the presence of a charge at the origin with semilength $\ell=1$ and any other charges are given solely by the probability of finding $q_{\pm\ell^{\prime}}^{\ell^{\prime}}$. Therefore, we find $\langle q_{m}^{1}q_{n}\rangle\sim|m-n|^{-3/2}$.
@@
For a charge to be relevant to thermodynamics, it must be extensive and additive; despite having power-law tails, the matched parentheses charge indeed exhibits both of these properties. To see this, consider a region of $2L\rightarrow\infty$ total sites split into two $L$-site subsystems $A$ and $B$. The total charge in this region $Q_{AB}=\sum_{n=1}^{2L}q_{n}$. The total number of matched parentheses charges in subsystem $A$, given by $Q_{A}=\sum_{n=1}^{L}q_{n}$, grows with $L$, and so $\langle Q_{A}\rangle\sim L$. One could worry that in spite of the extensivity of the matched parentheses charge, the power-law tails of the charge may lead to $Q_{AB}\neq Q_{A}+Q_{B}$. From showing exactly that $\langle Q_{A}^{2}\rangle_{c}\sim L$, we conclude that fluctuations of the total charge are sufficiently suppressed to guarantee additivity $Q_{AB}\simeq Q_{A}+Q_{B}$, up to subleading corrections of $\mathcal{O}(\sqrt{L})$ – for further details, see Appendix A. Therefore, fluctuations in the value of $Q_{A}$ are suppressed as $\sqrt{\langle Q_{A}^{2}\rangle_{c}}/{\langle Q_{A}\rangle}\sim 1/\sqrt{L}$, so the local density of matched parentheses is well-defined. Since the matched parentheses charge is locally conserved, extensive, and additive, it is relevant to hydrodynamics. Intuitively, due to the fact that each matched parentheses charge is tied to two individual parentheses, both of which move with $z\approx 5/2$, the matched parentheses charge should have the same dynamical exponent as the original charge– this intuition can be confirmed numerically (see Appendix C).